OpenStudy (anonymous):

Show that the equation has a root in the given interval. F(x)=-x^4+3x^2+5=0 What the heck are they asking me to do here???

OpenStudy (anonymous):

whts the interval?

OpenStudy (anonymous):

integrate this eqn

OpenStudy (anonymous):

then put inh both the boundary values of the interval [a,b] the values obtained will be same for a and b

OpenStudy (anonymous):

oh sorry, interval is 2,3

OpenStudy (anonymous):

so integrate ur eqwn and then put in 2 and 3..ull c the values obtained at 2 and 3 are equal hence proved

OpenStudy (anonymous):

so I am using 2 and just plugging it in, and then do the same with 3?

OpenStudy (anonymous):

after integrating the eqn..yes

myininaya (myininaya):

does F symbolize the the antiderivative

OpenStudy (anonymous):

@ssnapier: Are you in a calculus class?

OpenStudy (anonymous):

not necessary@myininaya

OpenStudy (anonymous):

this is for college algebra

myininaya (myininaya):

if F symbolizes the function we are trying to find a root, we just evaluate F at 2 and evaluate F at 3 if one is positive and one is negative, then we have a root on that interval

OpenStudy (anonymous):

Ah ok. Ignore the part about integrals then. That's like 2 years ahead of what we're dealing with. :P

OpenStudy (anonymous):

yeah...we can do that as well....i ws applying the rolles thm

myininaya (myininaya):

to apply rolle's thm the continous F must satify F(2)=F(3), then do F'(c)=0 and solve for c

OpenStudy (anonymous):

@him: you have to remember he's in college algebra. He hasn't even begun touching limits let alone derivatives and integrals and antiderivatives. There's gotta be an alternate way to do this (I just can't remember it).

myininaya (myininaya):

all you do is do F(3) and F(2) ;lol

OpenStudy (anonymous):

I am doing it right now...

OpenStudy (anonymous):

ummmm, I must have done this wrong...

OpenStudy (anonymous):

@uber: how cn i guess wht level hes studying @ssnapier : sorry mate

OpenStudy (anonymous):

I plugged in 2, and got 16+12+5=0 how can that be?? with 3 I got 81+27+5=0

myininaya (myininaya):

F(2)=1 F(3)=-49

OpenStudy (anonymous):

uve got a minus at the beginning i suppose..yeah shes right..

OpenStudy (anonymous):

holy hell... how did I ge that far off??

OpenStudy (anonymous):

but -2^4 is a positive 16, right?

OpenStudy (anonymous):

u ignored the minus sign at the beginning

OpenStudy (anonymous):

its -(x^4)

myininaya (myininaya):

the negative is not being raised just the x

OpenStudy (anonymous):

ahhhhh, that makes much more sense...

myininaya (myininaya):

since F(2)>0 and F(3)<0, then there is a number between 2 and 3 lets call it c such that F(c)=0

OpenStudy (anonymous):

bingo

OpenStudy (anonymous):

I can't believe that such a simple concept actually was accredited to someone. lulz.

myininaya (myininaya):

maybe it wasn't so easy to see back then

OpenStudy (anonymous):

Then again... I disagree with the theorem. What if there's a hole at where y would equal zero? I mean I know it doesn't have a hole here because we've got a polynomial, but what if it was something like "in the interval 3,5... prove that -blahblah/(x-4) has a zero" At 4, it'd die. And I'd make the equation purposely have a limit approaching zero at 4 :P

myininaya (myininaya):

the function has to be continuous

myininaya (myininaya):

intermediate value theroem

OpenStudy (anonymous):

Then I'll fall back to my previous point. :P As long as Cartesius hadn't invented his graphing system yet, there's no reason for anyone to be surprised by this logic. But if it was before the graph was invented, then this guy deserves his credits. :D

myininaya (myininaya):

to apply any theorems it must also satisfy the hypothesis of that theorem